OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 3x + 7y = 11? please help i will give medals too :) <3

OpenStudy (anonymous):

You have one equation and 2 variables. You will have infinite number of solutions. Pick any number for x, solve for y, and that's the ordered pair.

OpenStudy (anonymous):

a(1, –2) B.(0,–4) c.(6, –1) d.(0, 4)

OpenStudy (anonymous):

If you have pairs to choose from, just substitute them and see which one works. for: a) 3*(1)+7*(-2)=11 Is this true? b) 3*(0)+7*(-4)=11 Is this true? c) 3*(6)+7*(-1)=11 Is this true? d) 3*(0)+7*(4)=11 Is this true? The one that is true, is the ordered pair that is the solution.

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

This is a) Is 3 times 1 plus 7 times negative 2 equal to 11 ? If it is equal to 11, then (1,-2) is the solution to the question